xaguia2:

Which expression is equivalent to 5(2+7)? A. 2(5+7) B. 2+7(5) C. 5(2)+7 D.5(2)+5(7) @thisgirlpretty @shadow @563blackghost @angle

Angle:

Hello :) there are multiple ways to do this for example, if you were to not look at any of the options, what answer would you get from just 5*(2 + 7) = ?

Angle:

if you're not comfortable with the distributive property yet, you can do 5*(2 + 7) = 5 * 9 = 45

Angle:

then we can simplify the rest of the options and see which one equals 45

jasonmitchell:

Just distribute. 5(2 + 7) 10 + 35 45